# Holomorphic Yang-Mills fields on $B$-branes

### v1

@@ Abstract.
Considering the $B$-branes over a complex manifold $Y$ as objects of the bounded derived category $D^{b}(Y)$, we define holomorphic gauge fields on $B$-branes and the Yang-Mills functional for these fields. These definitions are a generalization to $B$-branes of concepts that are well known in the context of vector bundles. Given ${\mathscr{F}}^{\bullet}\in D^{b}(Y),$ we show that the Atiyah class $a({\mathscr{F}}^{\bullet})\in{\rm Ext}^{1}({\mathscr{F}}^{\bullet},\,\Omega^{1}({\mathscr{F}}^{\bullet}))$ is the obstruction to the existence of gauge fields on ${\mathscr{F}}^{\bullet}.$ When $Y$ be either the projective space ${\mathbb{P}}^{n}$ or the variety of complete flags in ${\mathbb{C}}^{3},$ we prove that the cardinal of the set of holomorphic gauge fields on any $B$-brane over $Y$ is $\leq 1.$ We prove that the set of Yang-Mills fields on the $B$-brane ${\mathscr{F}}^{\bullet},$ if it is nonempty, is in bijective correspondence with the points of an algebraic subset of ${\mathbb{C}}^{m}$ defined by $m$ polynomial equations of degree $\leq 3$, where $m={\rm dim}\,{\rm Ext}^{0}({\mathscr{F}}^{\bullet},\,\Omega^{1}({\mathscr{F}}^{\bullet})).$
@@ 1. Introduction
In this article, we extend the well-known concepts of gauge field and Yang-Mills field on vector bundles to $B$-branes. From the mathematical point of view, a $B$-brane over a complex manifold $Y$ is an object of $D^{b}(Y)$, the bounded derived category of coherent analytic sheaves over $Y$ [2, Sect. 5.4] [3, Sect. 5.3].
@@
The simplest $B$-banes are the holomorphic vector bundles. A gauge field on a holomorphic vector bundle $V\to M,$ in mathematical terms a connection on $V,$ defines a derivation of sections along vectors of $V$ giving rise to sections not necessarily holomorphic. A holomorphic gauge field on $V,$ i., e. holomorphic connection on $V$ [4], allows us to define a derivative of the holomorphic sections of $V$ along any “direction”, giving rise to holomorphic sections.
@@
But no every holomorphic vector bundle admits a holomorphic connection. The obstruction for this existence is the Atiyah class of the bundle [4]. Hence, the vanishing of this class is a necessary condition for the existence of such gauge fields on the vector bundle. Assuming that there is no obstruction to the existence of holomorphic gauge fields, the set of these fields on a vector bundle over a compact manifold is a finite dimensional affine space, in contrast to the smooth case. This property is a consequence of the finite dimensionality of the coherent cohomology groups.
@@
The two features discussed above concerning the holomorphic connections on vector bundles: the possible obstruction to its existence, and the finite dimension of the space of those fields, when it is not empty, are also present in the extension of the holomorphic gauge fields to general branes.
Gauge fields on coherent sheaves. A gauge field on $V$ determines an isomorphism between the fibers over “close” points of $M$. In fact, the connection is determined by the identification of the fibers at points “infinitesimally close” in $M;$ i. e. by the identifications in the directions defined by the tangent vectors of the base. The idea of being “infinitesimally close” can be formulated by means of the first infinitesimal neighborhood of the diagonal of $M$. In this setting, a gauge field on $V$ can be regarded as a right inverse of the projection $\pi:J^{1}(V)\to V,$ where $J^{1}(V)$ is the $1$-jet bundle of $V$ [30].
That approach admits a natural translation to the context of the coherent sheaves over complex manifolds. In this way, if ${\mathscr{F}}$ is a coherent sheaf over the compact analytic manifold $Y,$ we define a holomorphic gauge field on ${\mathscr{F}}$ as a right inverse of the natural morphism ${\mathscr{J}}^{1}({\mathscr{F}})\to{\mathscr{F}}$, where ${\mathscr{J}}^{1}({\mathscr{F}})$ is the corresponding $1$-jet sheaf. This inverse determines a morphism of abelian sheaves
$\nabla:{\mathscr{F}}\to\Omega^{1}({\mathscr{F}})=\Omega^{1}_{Y}\otimes_{{\mathscr{O}}_{Y}}{\mathscr{F}},$
@@
which satisfies the Leibniz’s rule. In this context, the obstruction to the existence of a holomorphic connection on the sheaf ${\mathscr{F}}$ is an element $a({\mathscr{F}})$ of the group ${\rm Ext}^{1}({\mathscr{F}},\,\Omega^{1}({\mathscr{F}}))$. Furthermore, when the set of holomorphic gauge fields on ${\mathscr{F}}$ is nonempty, it is an affine space associated to the finite dimensional vector space ${\rm Hom}({\mathscr{F}},\,\Omega^{1}({\mathscr{F}}))$.
@@
Yang-Mills fields on shaves. Setting ${\mathcal{S}}$ for the singularity set of ${\mathscr{F}}$, let us assume that on the locally free sheaf ${\mathscr{F}}|_{Y\setminus{\mathcal{S}}}$ there is defined a Hermitian metric, we say that ${\mathscr{F}}$ is a Hermitian sheaf. If $Y$ is a Kähler manifold, on the set of holomorphic connections over ${\mathscr{F}}$, one defines the Yang-Mills functional $\mathcal{YM}$. Essentially, the value of $\mathcal{YM}$ at a connection $\nabla$ is the squared norm $\lVert K_{\nabla}\rVert^{2}$ of the curvature of that connection [14, page 417], [27, page 44].
The stationary points of functional $\mathcal{YM}$ are the Yang-Mills gauge fields on ${\mathscr{F}}$. The set of critical points of this functional will be denoted by ${\sf YM}({\mathscr{F}})$.
@@
The reflexive sheaves might be thought as “vector bundles with singularities” [16, page 121], and these singularities can in some way be avoided. From this property, and denoting $m:={\rm dim}\,{\rm Hom}({\mathscr{F}},\,\Omega^{1}({\mathscr{F}})),$ we will prove that if the reflexive sheaf ${\mathscr{F}}$ admits holomorphic gauge fields, then the set ${\sf YM}({\mathscr{F}})$ is in bijective correspondence with the points of an algebraic subset of ${\mathbb{C}}^{m}$ defined by $m$ algebraic equations of degree $\leq 3$ (Theorem 1). Thus, in general, the set of Yang-Mills fields on ${\mathscr{F}}$ will be a finite set.
@@
Yang-Mills fields on a $B$-brane. As we said, a $B$-brane on $Y$ is a complex $({\mathscr{F}}^{\bullet},\,\delta^{\bullet})$ of analytic coherent sheaves on $Y.$ According to the preceding paragraphs, it is reasonable to define a gauge field on this brane as an element of ${\rm Hom}_{D^{b}(Y)}\big{(}{\mathscr{F}}^{\bullet},\,{\mathscr{J}}^{1}({\mathscr{F}}^{\bullet})\big{)}$ which lifts the identity on ${\mathscr{F}}^{\bullet}.$
Let $X$ be either the projective space ${\mathbb{P}}^{n}$ or the variety of complete flags in ${\mathbb{C}}^{3}.$ We prove that the cardinal of the set of holomorphic gauge fields on any $B$-brane over $X$ is $\leq 1$ (Theorems 3 and 4). For the proof of that result we use the existence of finite families of objects in $D^{b}(X)$ which generate those categories. A well-known theorem of Beilinson asserts that the set $\{{\mathscr{O}}_{{\mathbb{P}}^{n}}(-n),\,{\mathscr{O}}_{{\mathbb{P}}^{n}}(-n+1),\dots,{\mathscr{O}}_{{\mathbb{P}}^{n}}(-1),\,{\mathscr{O}}_{{\mathbb{P}}^{n}}\}$ is a full exceptional family in $D^{b}({\mathbb{P}}^{n})$ [19, Sect 8.3]. However, for the proof of the above result we need to construct a more appropriate generator set. For the case when $X$ is the mentioned flag variety, we determine a family of generators $D^{b}(X)$ following an idea from [5].
If $Y$ is a Hodge manifold, then it is a projective smooth variety and by the GAGA correspondence the analytic coherent sheaves on $Y$ can be considered as algebraic ones. Moreover, the category $D^{b}(Y)$ is equivalent to the homotopy category $K^{b}(Y)_{\rm coh}$ of complexes of injective sheaves with coherent and bounded cohomology. In this way, a gauge field $\psi$ on ${\mathscr{F}}^{\bullet}$ can be considered as a homotopy class of morphisms between the complexes of abelian sheaves ${\mathscr{F}}^{\bullet}\to\widehat{\Omega^{1}({\mathscr{F}}^{\bullet})}$, where $\widehat{\Omega^{1}({\mathscr{F}}^{\bullet})}$ is a complex in $K^{b}(Y)_{\rm coh}$ quasi-isomorphic to $\Omega^{1}({\mathscr{F}}^{\bullet})$. Thus, that homotopy class determines a unique morphism between each of the cohomology sheaves $\vartheta^{j}:{\mathscr{H}}^{j}({\mathscr{F}}^{\bullet})\to{\mathscr{H}}^{j}(\Omega^{1}({\mathscr{F}}^{\bullet})).$ In summary, the gauge field on ${\mathscr{F}}^{\bullet}$ defines a family of connections $\vartheta^{j}$ on the cohomology sheaves ${\mathscr{H}}^{j}({\mathscr{F}}^{\bullet})$.
@@
When the cohomology sheaves ${\mathscr{H}}^{j}$ are Hermitian, we define the value of the Yang-Mills functional on the above gauge $\psi$ as $\sum_{i}(-1)^{i}\lVert K_{\vartheta_{i}}\rVert^{2}.$ Thus, the Yang-Mills functional is a kind of Euler characteristic of the gauge field. Obviously, this definition of the Yang-Mills functional on branes generalizes the one given for coherent sheaves.
@@
The gauge field $\psi$ is a Yang-Mills field if it is a stationary point of Yang-Mills functional. We will prove that if the $\vartheta^{j}$ are Yang-Mills fields on the ${\mathscr{H}}^{j}$, so is $\psi$ on the brane (Proposition 14). Theorem 6 is the converse to Proposition 14. In Theorem 5, we generalize the result given in Theorem 1, about the cardinal of the set of Yang-Mills fields on sheaves, to a general brane ${\mathscr{F}}^{\bullet}$.
The article is organized in two sections. In Section 2, we define the gauge fields on a coherent sheaf. In the first subsections of this section, we revise the definition of the $1$-jet sheaf of a coherent sheaf ${\mathscr{F}}$ on a complex manifold $Y,$ because, although this is well known in algebraic geometry, it is not so well known in the community of mathematical physicists. The Yang-Mills functional is introduced in Subsection 2.2, where Theorem 1 is proved.
In Section 3 are considered the gauge fields on a general $B$-brane. Subsection 3.1 defines these fields and discusses the obstruction to their existence. Subsection 3.2 is devoted to the study of gauge fields on $B$-branes over ${\mathbb{P}}^{n};$ in this subsection we prove the above Theorem 3 concerning the cardinal of the set of such fields. Theorem 4, about the cardinal of the set of fields on $B$-branes over the above flag variety, is proved in Subsection 3.3. We show, in Subsection 3.4, how a gauge field on a $B$-brane field gives rise to connections in cohomologies. Finally, we define the Yang-Mills functional on gauge fields over general $B$-branes, showing the reasons on which this definition is based. We also prove Theorems 5 and 6 and Proposition 14.
2. Gauge fields on coherent sheaves
2.1. Gauge fields on a sheaf.
@@ 2. Gauge fields on coherent sheaves
As explained in the Introduction, the existence of a gauge field on a coherent sheaf over a complex manifold $Y$ should define an isomorphism between the stalks of the sheaf at any “infinitesimally close” points of $Y.$ In case Y is an algebraic variety the neighborhoods in the Zariski topology are too large to express the idea of “infinitesimal proximity”. Instead by means of the first infinitesimal neighborhood $Y^{(1)}$ of $Y,$ [13, page 698], it is possible to formulate this idea in a suitable way also for algebraic varieties.
If $R$ is a ${\mathbb{C}}$-algebra, ${\rm Hom}({\rm Spec}\,R,\,X)$ is the set of points of complex algebraic variety $X$ with values in $R$. Two points $x,y$ are infinitesimally close if the morphism $(x,\,y):{\rm Spec}\,R\to X\times X$ factorizes through $X^{(1)}$ [7, page 6]. Therefore, following Deligne, one defines a gauge field on the coherent ${\mathscr{O}}_{X}$-module ${\mathscr{G}}$ as an element of
(2.1) ${\rm Hom}_{X^{(1)}}(\pi_{1}^{*}{\mathscr{G}},\,\pi_{2}^{*}{\mathscr{G}}),$
@@
which is the identity on $X$, where $\pi_{1},\pi_{2}:X^{(1)}\to X$ are the projections and $\pi_{i}^{*}$ the corresponding inverse image functor. By the adjunction isomorphism, a gauge field on ${\mathscr{G}}$ is a morphism ${\mathscr{G}}\to\pi_{1*}\pi^{*}_{2}{\mathscr{G}}$ which is a right inverse of the projection $\pi:\pi_{1*}\pi^{*}_{2}{\mathscr{G}}\to{\mathscr{G}}.$
@@
As $\pi_{1*}\pi^{*}_{2}{\mathscr{G}}$ is the first jet sheaf of the coherent sheaf ${\mathscr{G}}$, in the following paragraphs we review the definition of the jet sheaf and also that of first neighborhood of the diagonal.
2.1.1. Neighborhood of the diagonal

### v2

@@ Abstract.
Considering $B$-branes over a complex manifold $X$ as objects of the bounded derived category of coherent sheaves over $X$, we define holomorphic gauge fields on $B$-branes and introduce the Yang-Mills functional for these fields. These definitions extend well-known concepts in the context of vector bundles to the setting of $B$-branes.
For a given $B$-brane, we show that its Atiyah class is the obstruction to the existence of gauge fields. When $X$ is the variety of complete flags in a $3$-dimensional complex vector space, we prove that any $B$-brane over $X$ admits at most one holomorphic gauge field.
Furthermore, we establish that the set of Yang-Mills fields on a given $B$-brane, if nonempty, is in bijective correspondence with the points of an algebraic set defined by $m$ complex polynomials of degree less than four in $m$ indeterminates, where $m$ is the dimension of the space of morphisms from the brane to its tensor product with the sheaf of holomorphic one-forms.
@@ 1. Introduction
In this article, we extend the well-known concepts of gauge field and Yang-Mills field on vector bundles to $B$-branes. From the mathematical point of view, a $B$-brane over a complex manifold $X$ is an object of $D^{b}(X)$, the bounded derived category of coherent analytic sheaves over $X$ [2, Sect. 5.4] [3, Sect. 5.3].
@@
The simplest $B$-branes are the holomorphic vector bundles. A gauge field on a holomorphic vector bundle $V\to M,$ in mathematical terms a connection on $V,$ defines a derivation of sections of $V$ along vector fields on $M$ giving rise to sections not necessarily holomorphic. A holomorphic gauge field on $V,$ i. e. a holomorphic connection on $V$ [4], allows us to define a derivative of the holomorphic sections of $V$ along any “direction”, giving rise to holomorphic sections.
@@
Not every holomorphic vector bundle admits a holomorphic connection. The obstruction for this existence is the Atiyah class of the bundle. The vanishing of this class is a necessary and sufficient condition for the existence of such connections on the vector bundle. Assuming that there is no obstruction to the existence of holomorphic connections, the set of these connections on a vector bundle over a compact manifold is a finite dimensional affine space, in contrast to the smooth case. This property is a consequence of the finite dimensionality of the coherent cohomology groups.
@@
The two features discussed above concerning the holomorphic connections on vector bundles: the obstruction to its existence, and the finite dimension of the space of those fields, when it is not empty, are also present in the extension to general branes.
Connections on coherent sheaves. Given a coherent sheaf ${\mathscr{F}}$ over the compact analytic manifold $X,$ a holomorphic connection on ${\mathscr{F}}$ [17] determines isomorphisms between the stalks of ${\mathscr{F}}$ over “close” points of $X;$ i. e. identifications in the directions defined by the tangent vectors of the base.
The idea of being “infinitesimally close” can be formulated by means of the first infinitesimal neighborhood of the diagonal of $X$. In this setting, a connection on ${\mathscr{F}}$ can be regarded as a right inverse of the natural morphism ${\mathscr{J}}^{1}({\mathscr{F}})\to{\mathscr{F}}$, where ${\mathscr{J}}^{1}({\mathscr{F}})$ is the corresponding $1$-jet sheaf [26, Sect. 3]. This inverse determines a morphism of abelian sheaves $\nabla:{\mathscr{F}}\to\Omega^{1}({\mathscr{F}})=\Omega^{1}_{X}\otimes_{{\mathscr{O}}_{X}}{\mathscr{F}},$ which satisfies the Leibniz rule.
@@
In this context, the obstruction to the existence of a holomorphic connection on the sheaf ${\mathscr{F}}$ is an element of the group ${\rm Ext}^{1}({\mathscr{F}},\,\Omega^{1}({\mathscr{F}})).$ Furthermore, when the set of holomorphic connections on ${\mathscr{F}}$ is nonempty, it is an affine space associated to the finite dimensional vector space ${\rm Hom}_{{\mathscr{O}}_{X}}({\mathscr{F}},\,\Omega^{1}({\mathscr{F}}))$.
@@
Yang-Mills fields on sheaves. The reflexive sheaves might be thought as “vector bundles with singularities” [13, p. 121], and these singularities can in some way be avoided. Setting ${\mathcal{S}}$ for the singularity set of the reflexive sheaf ${\mathscr{F}}$, let us assume that on the locally free sheaf ${\mathscr{F}}|_{X\setminus{\mathcal{S}}}$ there is defined a Hermitian metric, we say that ${\mathscr{F}}$ is a Hermitian sheaf. If $X$ is a Kähler manifold, on the set of holomorphic connections over ${\mathscr{F}},$ one defines the Yang-Mills functional $\mathcal{YM}$. Essentially, the value of $\mathcal{YM}$ at a connection $\nabla$ is the squared norm $\lVert K_{\nabla}\rVert^{2}$ of the curvature of that connection [12, p. 417], [22, p. 44], [23, p. 357].
@@
The stationary points of $\mathcal{YM}$ are the Yang-Mills fields on ${\mathscr{F}}$. The set of these points of this functional will be denoted by ${\sf YM}({\mathscr{F}})$. Denoting $m:={\rm dim}\,{\rm Hom}_{{\mathscr{O}}_{X}}({\mathscr{F}},\,\Omega^{1}({\mathscr{F}})),$ we will prove that if the reflexive sheaf ${\mathscr{F}}$ admits holomorphic connections, then the set ${\sf YM}({\mathscr{F}})$ is in bijective correspondence with the points of an algebraic subset of ${\mathbb{C}}^{m}$ defined by $m$ algebraic equations of degree $\leq 3$ (Theorem 3). Thus, in general, the set of Yang-Mills fields on ${\mathscr{F}}$ will be a finite set.
@@
Yang-Mills fields on a $B$-brane. As we said, a $B$-brane on $X$ is a complex $({\mathscr{F}}^{\bullet},\,\delta^{\bullet})$ of analytic coherent sheaves on $X.$ According to the preceding paragraph, it is reasonable to define a holomorphic gauge field on this brane as an element $\psi\in{\rm Hom}_{D^{b}(X)}\big{(}{\mathscr{F}}^{\bullet},\,{\mathscr{J}}^{1}({\mathscr{F}}^{\bullet})\big{)}$ which lifts the identity on ${\mathscr{F}}^{\bullet}.$
When the manifold $X$ admits a stratification, where the strata satisfy certain properties, the space of holomorphic gauge fields on any $B$-brane on $X$ is a set with cardinal $\leq 1$ (Theorem 1). The case where $X$ is the projective space ${\mathbb{P}}^{n}$ has been considered in [27]. Here, we study the case where $X$ is the variety of complete flags in ${\mathbb{C}}^{3}$ (Theorem 2).
The homomorphic gauge field $\psi$ determines a unique morphism between each of the cohomology sheaves $\vartheta^{j}:{\mathscr{H}}^{j}({\mathscr{F}}^{\bullet})\to{\mathscr{H}}^{j}(\Omega^{1}({\mathscr{F}}^{\bullet})),$ which in fact is a holomorphic connection ${\mathscr{H}}^{j}({\mathscr{F}}^{\bullet})$.
@@
When the cohomology sheaves ${\mathscr{H}}^{j}$ are Hermitian, we define the value of the Yang-Mills functional on the above gauge field $\psi$ as $\sum_{i}(-1)^{i}\lVert K_{\vartheta_{i}}\rVert^{2}.$ Thus, the Yang-Mills functional is a kind of Euler characteristic of the gauge field. Obviously, this definition of the Yang-Mills functional in the context of $B$-branes generalizes the one for coherent sheaves.
@@
The gauge field $\psi$ is a Yang-Mills field if it is a stationary point of Yang-Mills functional. In Theorem 4, we generalize the result given in Theorem 3, about the cardinal of the set of Yang-Mills fields on sheaves, to a general brane ${\mathscr{F}}^{\bullet}$. We will prove that if the $\vartheta^{j}$ are Yang-Mills fields on the ${\mathscr{H}}^{j}$, so is $\psi$ on the brane (Proposition 11). Theorem 5 is a partial converse to Proposition 11.
The article is organized in two sections. In Section 2 are considered the holomorphic gauge fields on $B$-branes. Subsection 2.1 concerns with the holomorphic connections on a sheaf. In Subsection 2.2, we define the holomorphic gauge fields on a brane and discuss the obstruction to their existence. Subsection 2.3 discusses branes on a manifold which admits certain stratifications; in this subsection we prove Theorems 1 and 2.
Section 3 is devoted to the holomophic Yang-Mills fields. In Subsection 3.1, we introduce the Yang-Mills functional for holomorphic connections on reflexive sheaves and prove Theorem 3. The Yang-Mills functional for holomorphic gauge fields on a $B$-brane is defined in Subsection 3.2, along with a justification for this definition. In this subsection are also proved Proposition 11 and Theorems 4 and 5.
2. Holomorphic gauge fields on $B$-branes
As we mentioned in the Introduction, our purpose is to define holomorphic gauge fields on $B$-branes, extending the concept of holomorphic connection on vector bundles.
2.1. Holomorphic connections on a sheaf.
@@ 2. Holomorphic gauge fields on $B$-branes
As explained in the Introduction, given a coherent sheaf ${\mathscr{F}}$ over the complex manifold $X$, the existence of a homomorphic connection on ${\mathscr{F}}$ should define an isomorphism between the stalks of the sheaf at any two “infinitesimally close” points of $X.$ By means of the first infinitesimal neighborhood $X^{(1)}$ of diagonal of $X$ [11, p. 698], it is possible to formulate this idea in a suitable way also for algebraic varieties.
The outline of a new definition for a connection on a sheaf ${\mathscr{G}}$ over a complex variety $X$, inspired by an idea from crystalline cohomology, is as follows. If $R$ is a ${\mathbb{C}}$-algebra, ${\rm Hom}({\rm Spec}\,R,\,X)$ is the set of points of complex algebraic variety $X$ with values in $R$. Two points $x,y$ are infinitesimally close if the morphism $(x,\,y):{\rm Spec}\,R\to X\times X$ factorizes through the infinitesimal neighborhood $X^{(1)}.$ Hence, a connection on the sheaf ${\mathscr{G}}$ should define an identification of the pullbacks $x^{*}{\mathscr{G}}$ and $y^{*}{\mathscr{G}}$ for any two infinitesimally close points.
Therefore, following Deligne [7, p. 6], one defines a connection on the coherent ${\mathscr{O}}_{X}$-module ${\mathscr{F}}$ as an element of
(2.1) ${\rm Hom}_{X^{(1)}}(\pi_{1}^{*}{\mathscr{F}},\,\pi_{2}^{*}{\mathscr{F}}),$
@@
which is the identity on $X$, where $\pi_{1},\pi_{2}:X^{(1)}\to X$ are the projections and $\pi_{i}^{*}$ the corresponding inverse image functor. By the adjunction isomorphism, a holomorphic connection on ${\mathscr{F}}$ is a morphism ${\mathscr{F}}\to\pi_{1*}\pi^{*}_{2}{\mathscr{F}}$ which is a right inverse of the projection $\pi:\pi_{1*}\pi^{*}_{2}{\mathscr{F}}\to{\mathscr{F}}.$
@@
As $\pi_{1*}\pi^{*}_{2}{\mathscr{F}}$ is the first jet sheaf of the coherent sheaf ${\mathscr{F}}$, in the following paragraph we review the definition of the jet sheaf.
2.1.1. The first jet sheaf